# Multimarked Spatial Search by Continuous-Time Quantum Walk

### 05 Jan 2024 v2

@@ 4 Final remarks
We have developed a framework to assess analytically the computational complexity of spatial search algorithms using continuous-time quantum walks on graphs with multiple marked vertices. This framework is underpinned by four propositions that not only lay the mathematical foundation but also offer a practical method for computing the eigenvalues and eigenvectors of the Hamiltonian. Additionally, these propositions facilitate the determination of the optimal value for the parameter $\gamma$ in the search algorithm. A key step in this process involves calculating a basis of eigenvectors of the kernel of a symmetric matrix, the dimension of which equals the number of marked vertices. This method applies to any graph whose associated Hamiltonian exhibits a sufficiently small gap between $\lambda^{\pm}$. In simpler scenarios, the success probability of the search algorithm can be represented as $a^{2}\sin^{2}\epsilon t$, with $2\epsilon$ usually being the spectral gap. The optimal running time is $t=\pi/2\epsilon$, leading to a success probability $p_{\mathrm{succ}}=a^{2}$. For the search algorithm to achieve optimality, implying a time complexity of $O(\sqrt{N})$, it is crucial that $\epsilon$ scales as $O(1/\sqrt{N})$ and $a$ maintains an order of $\Omega(1)$.
@@
The results of this work can accelerate numerical calculations of the computational complexity of search algorithms on arbitrary graphs. In this context, certain analytical expressions from Sections 2.2 and 2.3 can be circumvented. The numerical approach is outlined as follows. Start by determining numerically $\phi_{0}$ and $\big{|}\phi_{0}\big{\rangle}$. Next, use the entries of matrix $M^{\lambda}$, as defined in Eq. (3), to identify $\lambda^{\pm}$ and the optimal $\gamma$. Since $\lambda^{\pm}$ and $\gamma$ need to be found concurrently, we initially assign a small numerical value to $\gamma$, say $1/N^{2}$, and set $\lambda$ to $-\gamma\phi_{0}$. We then check whether det$(M^{\lambda})=0$. If this is not the case, we reduce $\lambda$ and continue the process. Upon locating the root, continue decreasing $\lambda$ to discover another root. Here, it becomes evident that $-\gamma\phi_{0}$ is not the midpoint. Consequently, increment $\gamma$ (and repeat the process of reducing $\lambda$) until the values that comply with the equation in Proposition 4 are found. Following this, $\epsilon$ is deduced from $|\lambda^{-}+\gamma\phi_{0}|$. We then proceed to numerically calculate $\big{\langle}\lambda^{\pm}\big{|}\phi_{0}\big{\rangle}$ and $\big{\langle}w\big{|}\lambda^{\pm}\big{\rangle}$ for all $w$. $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}$ and $\big{\langle}w\big{|}\lambda^{+}\big{\rangle}$ must be nonzero for at least one $w$, otherwise $\lambda^{+}$ does not fulfill the required conditions. Afterwards, we compute the success probability $p(t)$ using Eq. (11). By plotting $p(t)$, we can evaluate whether the curve resembles the square of a sinusoidal function and whether the peak is prominent.

### 17 Jan 2024 v3

@@ 4 Final remarks
We have developed a framework to assess analytically the computational complexity of spatial search algorithms using continuous-time quantum walks on graphs with multiple marked vertices. This framework is underpinned by four propositions that not only lay the mathematical foundation but also offer a practical method for computing the eigenvalues and eigenvectors of the Hamiltonian. Additionally, these propositions facilitate the determination of the optimal value for the parameter $\gamma$ in the search algorithm. A key step in this process involves calculating a basis of eigenvectors of the kernel of a symmetric matrix, the dimension of which equals the number of marked vertices. This method applies to any graph whose associated Hamiltonian exhibits a sufficiently small gap between $\lambda^{\pm}$. In simpler scenarios, the success probability of the search algorithm can be represented as $a^{2}\sin^{2}\epsilon t$, with $2\epsilon$ usually being the gap. The optimal running time is $t=\pi/2\epsilon$, leading to a success probability $p_{\mathrm{succ}}=a^{2}$. For the search algorithm to achieve optimality, implying a time complexity of $O(\sqrt{N})$, it is crucial that $\epsilon$ scales as $O(1/\sqrt{N})$ and $a$ maintains an order of $\Omega(1)$.
@@
The results of this work can accelerate numerical calculations of the computational complexity of search algorithms on arbitrary graphs. In this context, certain analytical expressions from Sections 2.2 and 2.3 can be circumvented. The numerical approach is outlined as follows. Start by determining numerically $\phi_{0}$ and $\big{|}\phi_{0}\big{\rangle}$. Next, use the entries of matrix $M^{\lambda}$, as defined in Eq. (3), to identify $\lambda^{\pm}$ and $\gamma$. Since they need to be found concurrently, we assign numerical values to both $\gamma$ and $\lambda$ and then check whether det$(M^{\lambda})=0$. If this is not the case, we change $\gamma$ and $\lambda$ until locating the roots. Following this, $\epsilon$ is deduced from $|\lambda^{-}+\gamma\phi_{0}|$. We then we proceed to numerically calculate $\big{\langle}\lambda^{\pm}\big{|}\phi_{0}\big{\rangle}$ and $\big{\langle}w\big{|}\lambda^{\pm}\big{\rangle}$ for all $w$. $\big{\langle}\lambda^{+}\big{|}\phi_{0}\big{\rangle}$ and $\big{\langle}w\big{|}\lambda^{+}\big{\rangle}$ must be nonzero for at least one $w$, otherwise $\lambda^{+}$ does not fulfill the required conditions. Afterwards, we compute the success probability $p(t)$ using Eq. (11). By plotting $p(t)$, we can evaluate whether the curve resembles the square of a sinusoidal function and whether the peak is prominent.